# Urban Visibility Hotspots: Quantifying Building Vertex Visibility from Connected Vehicle Trajectories using Spatial Indexing

### v1

@@ Abstract
Effective placement of Out-of-Home advertising and street furniture requires accurate identification of locations offering maximum visual exposure to target audiences, particularly vehicular traffic. Traditional site selection methods often rely on static traffic counts or subjective assessments. This research introduces a data-driven methodology to objectively quantify location visibility by analyzing large-scale connected vehicle trajectory data (sourced from Compass IoT) within urban environments. We model the dynamic driver field-of-view using a forward-projected visibility area for each vehicle position derived from interpolated trajectories. By integrating this with building vertex locations extracted from OpenStreetMap, we quantify the cumulative visual exposure, or “visibility count”, for thousands of potential points of interest near roadways. The analysis reveals that visibility is highly concentrated, identifying specific “visual hotspots” that receive disproportionately high exposure compared to average locations. The core technical contribution involves the construction of a BallTree spatial index over building vertices. This enables highly efficient (O(logN) complexity) radius queries to determine which vertices fall within the viewing circles of millions of trajectory points across numerous trips, significantly outperforming brute-force geometric checks. Analysis reveals two key findings: 1) Visibility is highly concentrated, identifying distinct ’visual hotspots’ receiving disproportionately high exposure compared to average locations. 2) The aggregated visibility counts across vertices conform to a Log-Normal distribution.
Index Terms:
Connected Vehicle Trajectories, Visibility Hotspots, Visibility Analysis, Spatial Indexing, Urban Analytics
I Introduction
@@ II Related Works
Traffic simulation including macroscopic, microscopic, and agent-based models (ABM), is a well-established tool in transportation research for analyzing network performance, safety, and demand [3, 4, 5]. While microscopic and ABM simulations offer the potential to model individual viewer (pedestrian or vehicle) paths and incorporate environmental factors affecting visibility [3, 6], their documented application specifically for quantifying geometric OOH visibility or optimizing placement based on such metrics appears limited. Approaches like Cellular Automata (CA) have been used to model the reach of mobile transit advertising, focusing on information spread rather than precise geometric visibility [7].
@@
Our work contributes to this landscape by: 1) Utilizing high-frequency (up to every 5 second trajectory updates) connected vehicle location and direction data which may offer broader coverage of traffic (includes passenger vehicles, buses, than specific sources like taxis or smart cards. 2) Focusing specifically on quantifying the geometric visibility of discrete urban features (building vertices from OSM) rather than relying solely on proximity or traffic volume metrics. 3) We also use BallTree spatial indexing explicitly for the efficient execution of numerous dynamic point-in-circle visibility queries derived from vehicle trajectories, enabling scalability. By combining these elements, we aim to provide a granular and computationally efficient method for assessing urban visibility potential from the perspective of real moving vehicles.
II-1 Study Area
@@
TABLE I: Geographical Bounding Box (WGS84 Coordinate System - EPSG:4326)
@@
Latitude $-33.91441^{\circ}$ S $-33.89325^{\circ}$ S
@@ III Methodology
III-A Spatial Querying and Visibility Calculation
@@
6:$N_{points}\leftarrow\textsc{GetNumberOfPoints}(B_{tree})$triangleright Total number of densified points
@@
11:         $(lat_{c},lon_{c})\leftarrow\textsc{CalculateDestinationPoint}(p.lat,p.lon,p.bearing,R_{meters})$triangleright Calculate circle center
@@
13:              $coord\_c^{rad}\leftarrow\textsc{ToRadians}([[lat_{c},lon_{c}]])$triangleright Convert center to radians
14:              $indices\leftarrow\textsc{QueryBallTreeRadius}(B_{tree},coord\_c^{rad},R_{angular})$triangleright Find points in circle
15:              for each index $idx$ in $indices[0]$ do $\triangleright$ Query returns list of lists
@@
24:         $p_{id}\leftarrow M_{id}[i]$triangleright Get original point ID from map
25:         $V_{trip}\{p_{id}\}\leftarrow C_{point\_counts}[i]$triangleright Store count in result map
@@ III-B Aggregation and Analysis
The primary goal here is to aggregate the visibility counts obtained for each densified building point during individual trips into a single total count. The aggregation logic proceeds as follows: 1) Per-trip visibility results (mapping point identifiers to counts) are processed, grouping counts by unique densified building point. A key derived from rounded coordinates is typically used to ensure these points are uniquely identified across different processing batches or potential floating-point discrepancies. 2) For each unique building point key, the visibility counts accumulated across all trips are summed to yield a final total_count. 3) The aggregated results are stored, typically in a structure mapping the unique point key back to its original coordinates and the calculated total_count. This aggregated count represents the overall frequency with which each densified building point along building outlines was found within a vehicle’s viewing area across the studied trajectories.
III-C Trajectory and Visibility Visualization
@@
III-D Aggregated Visibility Distribution
@@
III-E Statistical Distribution of Visibility
The aggregated visibility counts for the 34,495 unique building vertices with non-zero visibility were fitted to several common probability distributions. The goal was to characterize the overall distribution of visibility across the study area. Table III summarizes the estimated parameters for each fitted distribution, along with goodness-of-fit statistics including the Kolmogorov-Smirnov (K-S) statistic (D) and the 1-Wasserstein distance. Lower values for D and Wasserstein distance generally indicate a better fit to the empirical data.
TABLE II: Distribution Fitting Results for Aggregated Vertex Visibility Counts (N=34,495). Parameters shown are shape (s, a, c), location (loc), and scale. K-S p-values were effectively zero ($<10^{-70}$) for all fits shown.
@@ V Conclusion
VI ACKNOWLEDGEMENTS
We thank Compass IoT for the data and support provided for this study, especially to Mr David Lillo Trynes. This work has been funded by the UTS Jenny Edwards Fellowship awarded in 2025 to Assoc. Prof. Adriana-Simona Mihaita for conducting research on connected vehicles.

### v2

# Rapid Identification of Urban Visibility Hotspots: Quantifying Building Vertex Visibility from Connected Vehicle Trajectories using Spatial Indexing

@@ Abstract
Effective placement of Out-of-Home advertising and street furniture requires accurate identification of locations offering maximum visual exposure to target audiences, particularly vehicular traffic. Traditional site selection methods often rely on static traffic counts or subjective assessments. This research introduces a data-driven methodology to objectively quantify location visibility by analyzing large-scale connected vehicle trajectory data within urban environments. We model the dynamic driver field-of-view using a forward-projected visibility area for each vehicle position derived from interpolated trajectories. By integrating this with building vertex locations extracted from OpenStreetMap, we quantify the cumulative visual exposure, or “visibility count”, for thousands of potential points of interest near roadways. The core technical contribution involves the construction of a BallTree spatial index over building vertices. This enables highly efficient (O(logN) complexity) radius queries to determine which vertices fall within the viewing circles of millions of trajectory points across numerous trips, significantly outperforming brute-force geometric checks. Analysis reveals two key findings: 1) Visibility is highly concentrated, identifying distinct ’visual hotspots’ receiving disproportionately high exposure compared to average locations. 2) The aggregated visibility counts across vertices conform to a Log-Normal distribution.
Keywords: Connected Vehicle Trajectories, Visibility Hotspots, Visibility Analysis, Spatial Indexing, Urban Analytics.
fn1fn1footnotetext: ORCID: 0000-0001-6875-3568fn2fn2footnotetext: ORCID: 0000-0001-7670-5777
1 Introduction
@@ 2 Related Works
Traffic simulation including macroscopic, microscopic, and agent-based models (ABM), is a well-established tool in transportation research for analyzing network performance, safety, and demand [3, 4, 5]. While microscopic and ABM simulations offer the potential to model individual viewer (pedestrian or vehicle) paths and incorporate environmental factors affecting visibility [3, 6], their documented application specifically for quantifying geometric OOH visibility or optimizing placement based on such metrics appears limited. Approaches like Cellular Automata (CA) have been used to model the reach of mobile transit advertising, focusing on information spread rather than geometric shape of visibility [7].
@@
Our work contributes to the area by: 1) Utilizing high-frequency (up to every 5 second trajectory updates) connected vehicle location and direction data which may offer broader coverage of traffic (includes passenger vehicles, buses, than specific sources like taxis or smart cards. 2) Focusing specifically on quantifying the geometric visibility of discrete urban features (building vertices from OSM) rather than relying solely on proximity or traffic volume metrics. 3) We also use BallTree spatial indexing explicitly for the efficient execution of numerous dynamic point-in-circle visibility queries derived from vehicle trajectories, enabling scalability. By combining these elements, we aim to provide a granular and computationally efficient method for assessing urban visibility potential from the perspective of real moving vehicles.
2.1 Study Area
@@
Table 1: Geographical Bounding Box (WGS84 Coordinate System - EPSG:4326)
@@
Latitude $33.91441^{\circ}$ S $33.89325^{\circ}$ S
@@ 3 Methodology
Urban visibility assessment demands a geometric framework that balances analytical precision with computational scalability for large-scale transportation analytics. While vehicle forward-facing fields of view (FoVs) exhibit inherent angular constraints - typically 60-120 degrees in real-world driving scenarios - the primary objective here is not to model physical sensor limitations but to establish a computationally efficient baseline metric for visibility assessment.
This work adopts the circle as the foundational geometric approximation of driver FoV, explicitly prioritizing computational efficiency over precision. Processing $O(10^{6})$ vehicle trajectories with $O(10^{3})$ building points requires sub-second per-point operations to enable feasible large-scale analysis. The circle containment test ($O(1)$) executes in 6 arithmetic operations (2 subtractions, 2 multiplications, 1 addition, 1 comparison), while sector-based alternatives require $\sim$100-250 operations due to the $\arctan$ function (Table 2).
Table 2: Geometric containment algorithms and their computational cost
Shape & Algorithm Formula / Concept Computational Cost (Operations)
Circle
(Squared Distance Comparison)
Compares the squared distance from point $P(p_{x},p_{y})$ to center $C(c_{x},c_{y})$ with the squared radius $r^{2}$. This avoids a slow square root operation.
$(p_{x}-c_{x})^{2}+(p_{y}-c_{y})^{2}\leq r^{2}$
Fastest:
2 Subtractions
2 Multiplications
1 Addition
1 Comparison
(Total: 6)
Triangle
(Cross Product Test)
Checks if point $P$ is on the same side of all three edges (e.g., AB, BC, CA) by verifying the sign of the 2D cross product.
For edge AB: $z=(B_{x}-A_{x})(P_{y}-A_{y})-(B_{y}-A_{y})(P_{x}-A_{x})$
(Signs of all 3 results must be consistent)
Fast:
9 Subtractions
6 Multiplications
3 Comparisons
(Total: 18)
Sector
(Composite Check)
A two-part test:
1. Point must be inside the circle.
2. The angle of the point must be within the sector’s angular bounds.
$(p_{x}-c_{x})^{2}+(p_{y}-c_{y})^{2}\leq r^{2}\quad\land$
$\theta_{start}\leq\text{atan2}(p_{y}-c_{y},p_{x}-c_{x})\leq\theta_{end}$
Slow:
4 Subtractions
2 Multiplications
1 Addition
3 Comparisons
A computationally expensive ‘atan2‘ function (depends on the implementation).
Polygon ($n$-sided)
(Ray Casting Algorithm)
A Ray Casting Method counts how many times a ray extending from point $P$ intersects the polygon’s edges. An odd count means the point is inside.
$\text{is\_inside}=(\text{intersection\_count}\pmod{2}=1)$
(Requires a loop over all $n$ edges)
Slowest:
A loop of $n$ iterations. An optimized test may avoid slow division operation. The worst-case cost per iteration is:
4 Subtractions
2 Multiplications
5 Comparisons
(Total: $\approx$ 11 x n)
Selection of an Optimal Circular Sensor Approximation
For computational efficiency in real-time applications, it is often advantageous to approximate complex sensor geometries, such as a sector, with simpler shapes like a circle. A quantitative evaluation was performed to identify the optimal circular approximation for a sensor with a 100-meter range and a 60-degree field of view. Several candidate circles, defined by their maximum extent along the x-axis ($x_{\text{end}}$), were analyzed using a Monte Carlo simulation.
To provide a metric of similarity between the true sector (Area $A$) and each circular approximation (Area $B$), we utilize the Sørensen-Dice coefficient ($D$). This metric balances the penalties for both under-approximating (missed area) and over-approximating (excess area). It is defined as:
$D=\frac{2|A\cap B|}{|A|+|B|}$ (1)
A Dice coefficient of 1 indicates a perfect match, while 0 indicates no overlap. The results of this analysis, including the percentage of the sector’s area covered (Overlap), missed, and incorrectly included (Excess), are detailed in Table 3.
Table 3: Evaluation of circular approximations for a 60 degree visibility sector. The table shows the percentage of the sector’s area that is correctly covered (Overlap), missed (Missed), and incorrectly included (Excess). The Dice Coefficient measures the overall similarity, with the highest value indicating the best geometric fit.
$x_{\text{end}}$ (m) Overlap % Missed % Excess % Dice Coeff
80.00 58.63 41.47 37.51 0.60
90.00 74.15 25.95 47.48 0.67
95.00 82.57 17.52 52.88 0.70
100.00 91.47 8.63 58.57 0.73
105.00 96.78 3.32 68.57 0.73
110.00 99.28 0.82 82.12 0.71
115.47 100.10 0.00 99.71 0.67
We observe a clear trade-off: increasing the circle’s size reduces the Missed % at the cost of increasing the Excess %. While the circle with $x_{\text{end}}=115.47$ m guarantees complete coverage, which is critical for safety-first applications, it does so with substantial over-approximation, leading to a suboptimal Dice coefficient of 0.67. The approximation with $x_{\text{end}}=100.00$ m achieves the maximum Dice coefficient of 0.73. This indicates that it provides the best overall coverage, representing the most effective balance between capturing the true sensor area (91.5% overlap) and minimizing extraneous coverage. Consequently, for a balanced representation, the circle defined by $x_{\text{end}}=100.00$ m is selected as the optimal approximation. The visual comparison of these approximations is shown in Figure 2.
Figure 2: Visual comparison of circular approximations to the true sensor sector. The circle with $x_{\text{end}}=100.0$ m, highlighted with a solid line, provides the best overall similarity as measured by the Dice Coefficient
3.1 Spatial Querying and Visibility Calculation
@@
[…]
@@
11:   $(lat_{c},lon_{c})\leftarrow\textsc{CalculateDestinationPoint}(p.lat,p.lon,p.bearing,R_{meters})$ $\triangleright$ Calculate circle center
@@
13:     $coord\_c^{rad}\leftarrow\textsc{ToRadians}([[lat_{c},lon_{c}]])$ $\triangleright$ Convert center to radians
14:     $indices\leftarrow\textsc{QueryBallTreeRadius}(B_{tree},coord\_c^{rad},R_{angular})$ $\triangleright$ Find points in circle
15:     for each index idxidx in $indices[0]$ do $\triangleright$ Query returns list of lists
@@
24:   $p_{id}\leftarrow M_{id}[i]$ $\triangleright$ Get original point ID from map
25:   $V_{trip}\{p_{id}\}\leftarrow C_{point\_counts}[i]$ $\triangleright$ Store count in result map
@@ 3.2 Aggregation and Analysis
The primary goal here is to aggregate the visibility counts obtained for each densified building point during individual trips into a single total count. The aggregation logic proceeds as follows: 1) Per-trip visibility results (mapping point identifiers to counts) are processed, grouping counts by unique densified building point. A key derived from rounded coordinates is typically used to ensure these points are uniquely identified across different processing batches or potential floating-point discrepancies. 2) For each unique building point key, the visibility counts accumulated across all trips are summed to yield a final total_count. 3) The aggregated results are stored, in a structure mapping the unique point key back to its original coordinates and the calculated total_count. This aggregated count represents the overall frequency with which each densified building point along building outlines was found within a vehicle’s viewing area across the studied trajectories.
3.3 Trajectory and Visibility Visualization
@@
3.4 Aggregated Visibility Distribution
@@
3.5 Statistical Distribution of Visibility
The aggregated visibility counts for the 34,495 unique building vertices with non-zero visibility were fitted to several common probability distributions. The goal was to characterize the overall distribution of visibility across the study area. Table 5 summarizes the estimated parameters for each fitted distribution, along with goodness-of-fit statistics including the Kolmogorov-Smirnov (K-S) statistic (D) and the 1-Wasserstein distance. Lower values for D and Wasserstein distance generally indicate a better fit to the empirical data.
Table 4: Distribution Fitting Results for Aggregated Vertex Visibility Counts (N=34,495). Parameters shown are shape (s, a, c), location (loc), and scale. K-S p-values were effectively zero ($10^{-7}$) for all fits shown.
@@ 5 Conclusion
A notable limitation is the model’s inability to account for 3D occlusions. For example, at the intersection of Bourke Street and Ebsworth Street (coordinates: -33.905499, 151.203835), a tall commercial building blocks visibility to lower buildings despite high visibility counts. Future work will incorporate digital elevation models to address this constraint.
@@
6 ACKNOWLEDGEMENTS
We thank Compass IoT for the data and support provided for this study. This work has been funded by the UTS Jenny Edwards Fellowship awarded in 2025 to Assoc. Prof. Adriana-Simona Mihaita for conducting research on connected vehicles.